40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6255     Entry time: Fri Feb 3 23:19:22 2012
Author: Zach 
Type: Update 
Category: SUS 
Subject: OSEM testing begins 

I took one of the spare OSEM satellite amps (schematic) from the cabinet down the Y arm this afternoon to begin testing. I spent most of the day amassing the melange of adapters and connectors I needed to talk to the relic. The most elusive was the über-rare 64-pin IDE connector, for which neither the 40m nor Downs or Bridge had a breakout (despite there being several Phoenix boxes on each electronics rack at the 40m---hmm...). The solution I came up with was to make a breakout cable myself, only there was no 64-pin ribbon. So, I carefully fed a 50-pin and most of a 16-pin ribbon side by side into one push-down connector, and that was that:

IDE64_homestyle.png

I also finally found a 25-pin D-sub breakout just after figuring out the proper pinout for a 25-to-9 adapter, which I thought I was going to have to use. OH WELL.

Science

The first thing I figured I'd do is measure the LED drivers' current noise and see how it compared with LISO. I powered the box up and found that the TO-3 7815 regulator was putting out +20V---bad. I assumed it was broken, so I got another one from Downs and replaced it. Powered it up again and the output was still at +20V (WTF?). My suspicion is that one of the shielding capacitors has failed in some bizarre way, but I didn't have time to check this before I was beckoned to another task. This is where I'll start again next.

Another thing Frank and I noticed as we were figuring out how the driver worked was that the current-specifying resistor of one of the driver stages had not been properly modified along with the others, so it was forcing the feedback loop to rail. This mod was done precariously by adding two perpendicular sandwiched "Radd" resistors on top of the main one, so it's also possible that the ones for this stage had just been knocked off somehow (perhaps by the massive gender-switching ribbon chain hanging down on it). Steve and I noticed that there was a label on the box complaining that some part of the amp for one of the OSEMs wasn't working, but we peeled it off and threw it away because he figured it was outdated.

Anyway, in short, the plan going forward is as follows:

  • For this box
    • Measure the LED driver and PD transZ amp noises with dummy components
    • Compare with LISO to make sure they make sense
    • Measure the noise again with an OSEM connected
    • Based on the above and more LISO modeling, decide if it's a good idea to replace the LT1125's with OP497's
    • Increase the dynamic range by allowing +/-10V output, rather than +/-2V as was needed for old ADCs
  • After
    • Systematically mirror the changes in all other boxes by switching one out at a time

Comments welcome.

ELOG V3.1.3-